OpenStudy (anonymous):

A ball is thrown from an initial height of 3 ft with an initial upward velocity of 27ft/s The balls height h (in feet) after t seconds is given by the following. h=3+27t-16t^2. Find all values of t for which the balls height is 13 ft.

OpenStudy (anonymous):

we are looking for values of t for which h=13 -16t^2+27t+3=13 or -16t^2+27t-10=0 This is a quadratic equation with two solutions: t1=(27 - sqrt(89))/32 t2=(27 + sqrt(89))/32

OpenStudy (anonymous):

Thank you I knew it involved quadratic equations but couldn't remember what to do and seeing as I dont have my calculus book...